Solve using the four step process you learned.
A fenced-in rectangular area has a perimeter of 40 ft. The fence has a post every 5 ft.
a) How many posts are there?
b) What is the length and width of the field?
c) Are there any other possible answers to the above? If so, what?

Answer by stanbon(75887)

Solve using the four step process you learned.
A fenced-in rectangular area has a perimeter of 40 ft. The fence has a post every 5 ft.
Procedure::
1. Sketch a rectangle
2. Let the 2 shorter sides each be "x" in length
3. Then the 2 longer sides are (40-2x)/2 = 20-x in length
4. For the posts to be 5 ft apart the shorter sides must be x = 5 or x = 10
5. If x = 5, 20-x = 15 is the length of each longer side.
6. if x = 10, 20-x = 10 is the length of each longer side and the figure is a square.
7 Put a post at a corner and others 5 ft away from each successive post.
8 Count the posts.
---------
a) How many posts are there? 8
b) What is the length and width of the field? 10 by 10 or 5 by 15
c) Are there any other possible answers to the above? If so, what? no
------
Cheers,
Stan H.
--------